if f(x)= x^2+3, find f(5)
could you also show how you figured it out?

Answer by bucky(2189)

Given that f(x)= x^2+3
.
when you are asked to find f(5) that means that you are to find the value of the function when x equals 5.
.
So start with f(x)= x^2+3 and substitute 5 wherever you find an x to get:
.
f(5)= 5^2+3
.
square the 5 on the right side to get:
.
f(5) = 25 +3
.
and do the addition on the right side to get the answer:
.
f(5) = 28
.
Hope this helps you to understand what you are to do when you are asked to find f(x) in any similar problem. Just substitute the value of x that you are given into the right side of the equation and do the mathematics on the right side to get the answer.
.